OpenStudy (anonymous):

Which term is the perfect square of 3x4? 6x8 6x16 9x8 9x16

OpenStudy (anonymous):

@AlexandervonHumboldt2 @FLAMER123

Nnesha (nnesha):

square  take square of 3 and x^4

OpenStudy (alexandervonhumboldt2):

(3x4)^2=9x^8

OpenStudy (alexandervonhumboldt2):

(a^b)^c=a^bc

Nnesha (nnesha):

 x have power of 4 so multiply 4 by 2